OpenStudy (anonymous):

Find the derivative of arccos(x) through first principles. arccos(x) aka the inverse function of cos(x). How would you go about with doing this?

OpenStudy (anonymous):

first principles? are you not allowed to use the chain rule?

OpenStudy (anonymous):

OH YEAH!

OpenStudy (anonymous):

if you are allowed to use the chain rule and the fact that \[\cos(\arccos(x))=x\] then it is straighforward. if you have to use a limit good luck

OpenStudy (anonymous):

Wait how does that work? With the chain rule?

OpenStudy (anonymous):

in that case take the derivative of both sides and get \[\sin(\arccos(x))\times \arccos'(x)=1\] and then solve for \[\arccos'(x)\]

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

Wait, so you are starting with the property of how inverses can "undo" each other.

OpenStudy (anonymous):

you get \[\arccos'(x)=\frac{1}{-\sin(\arccos(x))}\]

OpenStudy (anonymous):

right and i made a mistake

OpenStudy (anonymous):

the derivative of cosine is -sine

OpenStudy (anonymous):

so we start with the identity \[\cos(\arccos(x))=x\] then take the derivative. you get \[-\sin(\arccos(x))\times \arccos'(x)=1\] by the chain rule (on the left)

OpenStudy (anonymous):

Ok wait, so you are doing: -sin(arccos(x))*arccox'(x) = 1

OpenStudy (anonymous):

we are pretending we don't know what \[\arccos'(x)\] is so we are going to solve for it using algebra

OpenStudy (anonymous):

Ok proceed please.

OpenStudy (anonymous):

first is it clear that the derivative of \[f(g(x))\] is \[f'(g(x))\times g'(x)\]?

OpenStudy (anonymous):

because that is what i used on the left hand side

OpenStudy (anonymous):

Yes so far.

OpenStudy (anonymous):

ok and in fact there is nothing special about cosine and arccosine. we could have just as well used f and f inverse

OpenStudy (anonymous):

so we are at this line \[-\sin(\arccos(x))\times \arccos'(x)=1\]

OpenStudy (anonymous):

i solve for \[\arccos'(x)\] in one simple algebra step and get \[\arccos'(x)=\frac{1}{-\sin(\arccos(x))}\]

OpenStudy (anonymous):

what the heck? Why is my solution to someone else's question up there?

OpenStudy (anonymous):

I just opened another window to answer someone else's question, and the answer I gave them is on this question..

OpenStudy (anonymous):

of course we are not done, but that is the idea. we now have to compute \[-\sin(\arccos(x))\] which is some trig work, but before we do it notice that if i replace cosine by f and arccosine by f inverse we get \[\frac{d}{dx}f^{-1}(x)=\frac{1}{f'(f^{-1}(x)}\]

OpenStudy (anonymous):

a very general statement that says if i want the derivative of an inverse function, i can find it by taking the reciprocal of the derivative of the original function evaluated at the inverse

OpenStudy (anonymous):

one minute. We started off with: cos(arccos(x)) = x right?

OpenStudy (anonymous):

again wrong! \[-\sin(\arccos(x))=-\sqrt{1-x^2}\]

OpenStudy (anonymous):

Who are you talking to?

OpenStudy (anonymous):

yes we started there. then take the derivative of both sides

OpenStudy (anonymous):

OHH Ok I see.

OpenStudy (anonymous):

the derivative of x is 1

OpenStudy (anonymous):

Got it. Ok then countinue from 1/-sin(arccos(x) = arccos'(x)

OpenStudy (anonymous):

the derivative of \[\cos(\arccos(x))\] is \[-\sin(\arccos(x))\times \arccos'(x)\] via the chain rule

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

so we have arccos'(x) = 1/-sin(arccos(x))

OpenStudy (anonymous):

right that line is correct

OpenStudy (anonymous):

now we need to find an algebraic expression for \[-\sin(\arccos(x))\] do you know how to get that one?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok lets call arcos(x) say theta, the angle whose cosine is x. then make a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

that ugly theta is supposed to represent arccos(x), the angle whose cosine is x

OpenStudy (anonymous):

that is why i labeled the "adjacent" side x and they hypotenuse 1

OpenStudy (anonymous):

because x / 1 = x

OpenStudy (anonymous):

now we want the sine of that angle, which amounts to finding the other side, which we do in one step by pythagoras. it is \[\sqrt{1-x^2}\]

OpenStudy (anonymous):